# Multiple importance sampling for stochastic gradient estimation

### v1

@@ 7 Limitations and future work
Our current OMIS implementation is limited to sequential models; hence its absence from our ViT experiment in Figure 5. However, there is no inherent limitation that would prevent its use with such more complex architectures. We anticipate that similar improvements could be achieved, but defer the exploration of this extension to future work.
@@
Recent work from Santiago et al. (2021) has explored a variant of importance sampling that forgoes sample-contribution normalization, i.e., the division by the probability in Equation 3 (and on line 10 of Algorithm 1). This heuristic approach lacks proof of convergence but can achieve practical improvement over importance sampling in some cases. We include a such variant of our IS method in Appendix F.
@@ 8 Conclusion
Ethic and impact
This paper presents work that aims to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Optimization details
The classification tasks include image classification (MNIST (Deng, 2012), CIFAR-10/100 (Krizhevsky et al., 2009) and point-cloud (ModelNet40 (Wu et al., 2015)) classification.
The MNIST database contains 60,000 training images and 10,000 testing images. We train a 3-layer fully-connected network (MLP) for MNIST over 50 epochs with an Adam optimizer (Kingma & Ba, 2014). CIFAR-10, introduced by Krizhevsky et al. (2009), is a dataset that consists of 60,000 color images of size 32x32. These images belong to 10 different object classes, each class having 6,000 images. On the other hand, CIFAR-100 Krizhevsky et al. (2009) contains 100 classes with 600 images each. For each class, there are 500 training images and 100 testing images. In our experiments, we train the ResNet-18 network (He et al., 2016) on both datasets. We apply random horizontal flip and random crops to augment the data during training. ModelNet40 contains 9,843 point clouds for training and 2,468 for testing. Each point cloud has 1,024 points. We train a PointNet (Qi et al., 2017) with 3 shared MLP layers and 2 fully-connected layers for 300 epochs on point-cloud classification. We use the Adam optimizer Kingma & Ba (2014), with batch size 64, weight decay 0.001, initial learning rate 0.00002 divided by 2 after 100, 200 epochs.
@@
Polynomial regression consists of optimizing the coefficients of a 1D polynomial of a given order to fit randomly drawn data from a reference polynomial of the same order. The reference data are generated on the interval . Optimization is done using an Adam optimizer (Kingma & Ba, 2014) with a mini-batch size of 32 elements.
The image regression task consists in learning the mapping between a 2D coordinate input (pixel coordinate) and the 3-color output of the image for this pixel. We use a network with 5 fully-connected layers associated with positional encodings using SIREN activations (Sitzmann et al., 2020). The training is done over 500 epoch using an Adam (Kingma & Ba, 2014) optimizer and each mini-batch is composed of pixels for a reference image.
@@ Appendix B Multiple importance sampling in brief
When dealing with multivariate functions, finding a probability density proportional to every parameters is often impractical. A trade-off is required to obtain a single probability distribution maximizing the proportionality with all the parameters of the function simultaneously. Several studies, such as (Zhao & Zhang, 2015; Needell et al., 2014; Wang et al., 2017; Alain et al., 2015), have shown that the optimal choice of sampling strategy is the norm of the function .
@@
The balance heuristic (Veach, 1997) is the most commonly used MIS heuristic. It sets the weight of the samples from each technique according to the following equation:
@@
Following the discussion in Section 5, it can also be deduced from Equations 11 and 6 that . Given a set of probability distribution functions , …, , we can formulate the optimal MIS solver as Algorithm 3. represents the vector containing the balance weight (13) w.r.t. the J sampling techniques and the normalization factor .
The algorithm proceeds through three key stages. The first stage involves initializing the linear system defined in Equation 7 (line 1). The second stage iteratively updates the system for each drawn data sample (lines 5-6). Upon completion of this process, the matrix and vector provide Monte Carlo approximations of the quantities specified in Equation 7. The third and final stage involve solving the linear system to obtain the vector (line 7). The estimated value of is then returned.
@@ Appendix F Additional results
This section provides additional results, including an ablation study as shown in Figure 9 for linear-system momentum used in Algorithm 2 and results of our adaptive sampling method. Figures 9 and 6 demonstrate that classical SGD, DLIS and Our IS work similarly with and without momentum. Our OMIS outperforms other methods in both cases.

### v2

@@ 7 Limitations and future work
Our current OMIS implementation is limited to sequential models; hence its absence from our ViT experiment in Fig. 5. However, there is no inherent limitation that would prevent its use with such more complex architectures. We anticipate that similar improvements could be achieved, but defer the exploration of this extension to future work.
@@
Recent work from Santiago et al., (2021) has explored a variant of importance sampling that forgoes sample-contribution normalization, i.e., the division by the probability in Eq. 3 (and on line 10 of Algorithm 1). This heuristic approach lacks proof of convergence but can achieve practical improvement over importance sampling in some cases. We include a such variant of our IS method in Appendix F.
@@ Appendix A Optimization details
The classification tasks include image classification (MNIST Deng, (2012), CIFAR-10/100 Krizhevsky et al., (2009) and point-cloud (ModelNet40 Wu et al., (2015)) classification.
The MNIST database contains 60,000 training images and 10,000 testing images. We train a 3-layer fully-connected network (MLP) for MNIST over 50 epochs with an Adam optimizer Kingma and Ba, (2014). CIFAR-10, introduced by Krizhevsky et al., (2009), is a dataset that consists of 60,000 color images of size 32x32. These images belong to 10 different object classes, each class having 6,000 images. On the other hand, CIFAR-100 Krizhevsky et al., (2009) contains 100 classes with 600 images each. For each class, there are 500 training images and 100 testing images. In our experiments, we train the ResNet-18 network He et al., (2016) on both datasets. We apply random horizontal flip and random crops to augment the data during training. ModelNet40 contains 9,843 point clouds for training and 2,468 for testing. Each point cloud has 1,024 points. We train a PointNet Qi et al., (2017) with 3 shared MLP layers and 2 fully-connected layers for 300 epochs on point-cloud classification. We use the Adam optimizer Kingma and Ba, (2014), with batch size 64, weight decay 0.001, initial learning rate 0.00002 divided by 2 after 100, 200 epochs.
@@
Polynomial regression consists of optimizing the coefficients of a 1D polynomial of a given order to fit randomly drawn data from a reference polynomial of the same order. The reference data are generated on the interval . Optimization is done using an Adam optimizer Kingma and Ba, (2014) with a mini-batch size of 32 elements.
The image regression task consists in learning the mapping between a 2D coordinate input (pixel coordinate) and the 3-color output of the image for this pixel. We use a network with 5 fully-connected layers associated with positional encodings using SIREN activations Sitzmann et al., (2020). The training is done over 500 epoch using an Adam Kingma and Ba, (2014) optimizer and each mini-batch is composed of pixels for a reference image.
@@ Appendix B Multiple importance sampling in brief
When dealing with multivariate functions, finding a probability density proportional to every parameters is often impractical. A trade-off is required to obtain a single probability distribution maximizing the proportionality with all the parameters of the function simultaneously. Several studies, such as Zhao and Zhang, (2015); Needell et al., (2014); Wang et al., (2017); Alain et al., (2015), have shown that the optimal choice of sampling strategy is the norm of the function .
@@
The balance heuristic Veach, (1997) is the most commonly used MIS heuristic. It sets the weight of the samples from each technique according to the following equation:
@@
Following the discussion in Section 5, it can also be deduced from Eqs. 11 and 6 that . Given a set of probability distribution functions , …, , we can formulate the optimal MIS solver as Algorithm 3. represents the vector containing the balance weight (13) w.r.t. the J sampling techniques and the normalization factor .
The algorithm proceeds through three key stages. The first stage involves initializing the linear system defined in Eq. 7 (line 1). The second stage iteratively updates the system for each drawn data sample (lines 5-6). Upon completion of this process, the matrix and vector provide Monte Carlo approximations of the quantities specified in Eq. 7. The third and final stage involve solving the linear system to obtain the vector (line 7). The estimated value of is then returned.
@@ Appendix F Additional results
This section provides additional results, including an ablation study as shown in Fig. 9 for linear-system momentum used in Algorithm 2 and results of our adaptive sampling method. Figs. 9 and 6 demonstrate that classical SGD, DLIS and Our IS work similarly with and without momentum. Our OMIS outperforms other methods in both cases.